Biplanes (56,11,2) with involutory automorphism fixing 14 points Vladimir Ćepulić and Mario Essert, Zagreb, Croatia

Keywords: biplane, involutory automorphism, automorphism group, orbital structure, indexing AMS Subject classification: 05B05 Abstract: A biplane (56,11,2) admitting an involutory automorphism fixing 14 points is isomorphic to one of the three known biplanes (56,11,2) with automorphism groups of order 80640, 288 and 64, respectively.